OpenStudy (anonymous):

Which of the following is not equivalent to the formula for the area of a rectangle A = LW? A over L= W A over W= L A = L over W A over WL= 1

OpenStudy (anonymous):

c.

OpenStudy (anonymous):

@texaschic101 @ganeshie8

OpenStudy (johnweldon1993):

\[\large A = LW\] Well if we divide both sides by W we get \[\large L = \frac{A}{W} \] or if we divided instead both sides by L we would get \[\large W = \frac{A}{L}\] and even if we want to divide both sides by LW we would then get \[\large \frac{A}{LW} = 1\] so which one is NOT listed here?

OpenStudy (anonymous):

C?

OpenStudy (anonymous):

@johnweldon1993 it's C right?

OpenStudy (anonymous):

@ganeshie8

OpenStudy (johnweldon1993):

Correct

OpenStudy (anonymous):

ok thx